# Large Language Model Guided Self-Debugging Code Generation

### v1

@@ 4.2. Results and Analysis
Table 1. Success rate comparison on three popular LLMs. We provide the mean value and standard deviation for PyCapsule with three experiment repeats. The results of other methods are from their official reports.
Method HumanEval HumanEval-ET MBPP MBPP-ET BigCodeBench
@@
AgentCoder [^1] Could not be reproduced during our experiments.  (Huang et al., 2024b) 96.3 86.0 91.8 91.8
@@
AgentCoder [^1] Could not be reproduced during our experiments.  (Huang et al., 2024b) 79.9 77.4 89.9 89.1
@@
Unlike other multi-agent architectures discussed in Section 2 that heavily rely on code generation agents, PyCapsule has more effective coordination among the computational overhead reduction, LLM’s API calls, and token processing.
In Table 1, PyCapsule achieves the state-of-the-art code generation tasks across HumanEval, HumanEval-ET, and BigCodeBench. While showing strong performance on MBPP, it falls short in comparison to AgentCoder and Qwen-2.5 (Hui et al., 2024) which employs three-shot prompt engineering to stabilize code generation. Although it shows remarkable potential, we maintain a consistent prompting approach across datasets to ensure comparability and utilise our signature converter module to generate actionable function signatures. Remarkably, on HumanEval, the integration of Qwen-2.5-Coder-Instruct’s 7B model with PyCapsule achieves a 94.1% success rate compared to 92.7% by its 32B model. Furthermore, on BigCodeBench, PyCapsule using Qwen-2.5-Coder-Instruct’s 7B shows a 25% improvement over the 49.6% reported by the 32B model variant. This demonstrates a significant savings in computational power while exceeding the performance.
Figure 2. Distribution of relative success ratios along the self-debugging attempts. The relative success ratio refers to the number of successful test cases with given attempts over the total number of successful test cases, measured by unit %. Results on all the datasets consistently show a decreasing relative success ratio as the number of attempts increases.
Influence of Self-debugging Attempts.
To evaluate the influence of each self-debugging attempt on the success rate, we analyze the normalized and accumulated impacts of each incremental attempt using GPT-4-1106-preview, GPT-3.5-Turbo-1106, and Qwen-2.5-Coder-Instruct models as the LLM for code generation. The success rate normalization was performed as follows: given $N$ problems, the number of successfully solved problems is denoted as $S_{0}$ while the rest $N_{1}=N-S_{0}$ unsolved problems will be further resolved with a self-debugging attempt. Accordingly, the number of problems solved with $i\geq 1$ attempts is denoted as $S_{i}$, and $N_{i+1}=N_{i}-S_{i}$ for the rest. The independent influence is defined as
@@
Figure 3. The left subfigure shows the normalized influence of each self-debugging attempt through independent accuracy improvements. The right subfigure highlights the mean influence of each self-debugging attempt on the overall accuracy through the accumulated influence without normalization.
This metric highlights the independent contribution of each attempt. In Figure 3, the descending trend in the independent influence indicates diminishing effectiveness of debugging attempts, where problems requiring more attempts may need enriched conversation history or enhanced prompts to improve the accuracy. We hypothesize that the first few self-debugging attempts address the most significant and easily identifiable errors while substantially improving the success rate. However, since these primary issues are resolved, more subtle and complex issues occur, and they are difficult to be detected and solved even with more attempts, resulting in diminishing normalised accuracy while increasing the overall accuracy by small amounts, as demonstrated by the right plot of Figure 3 . Additionally, the accumulation of minor fixes may introduce unseen issues that further degrade the effect of self-debugging attempts.
@@ 4.3. Discussion
PyCapsule’s single-attempt accuracy is significantly lower than Qwen-2.5’s reported performance, and it varies considerably across datasets, as shown in Figure 2 and detailed in Appendix A. Nevertheless, PyCapsule achieves higher single-attempt accuracy compared to GPT models implemented in MapCoder and AgentCoder. This nuanced performance difference highlights the varying impacts of prompt engineering and framework structure on code generation without self-debugging. With more debugging attempts, PyCapsule shows overall remarkable improvements, surpassing all the other self-debugging methods. This iterative improvement highlights the effectiveness of prompt engineering and error-handling ability of PyCapsule. Prior works (Wei et al., 2022; Chen et al., 2024; Patel et al., 2024) have demonstrated how prompt design significantly influences model performance, especially for generative AI tasks. PyCapsule leverages sophisticated system prompts that are dynamically augmented in the code generation and fix modes. Furthermore, an important advantage of PyCapsule is its reduced reliance on LLM-based agents, leading to reduced token consumption and improved computational efficiency. Existing frameworks like MapCoder can make up to 17 API calls with GPT models per problem on HumanEval and 12 calls on MBPP, following $k\times t$ debugging attempts, where $k$ is the number of retrieved exemplars and $t$ is the number of self-debugging attempts. In contrast, PyCapsule streamlines this process by limiting the programmer agent with a maximum of 5 attempts, requiring at most 6 LLM API calls for each problem because planning and code generation are handled within a single API call. Our empirical analysis shows that PyCapsule achieves superior code generation ability while maintaining efficient API usage, requiring an average of only 2.09, 1.38, and 1.55 API calls for HumanEval using GPT-3.5-Turbo-1106, GPT-4-Preview-1106, and Qwen-2.5-Coder-Instruct-7B, respectively. This optimization demonstrates the possibility of achieving high performance with significantly reduced computational overhead.
@@ Appendix A Contribution of Each Debugging Attempt
Appendix B HumanEval Experiments
@@
Appendix C Modules (Detailed)
C.1. Signature Converter
  • The inferred function name.
  • A structured function signature (e.g., argument names and types).
  • An example function call derived from the test case.
### Required function name for your reference ‘foo()’
### Function signature for your reference - foo(arg_int: int)
### An example function call from private test cases - foo(4)
C.2. Example Call Detection
C.3. Error-Handling
@@ Appendix D Pseudo-code of PyCapsule
Appendix E System Prompts
E.1. Code Generation Mode

### v2

@@ 4.2. Results and Analysis
Table 1. Success rate comparison on three popular LLMs. We provide the mean value and standard deviation for PyCapsule from three experiment repeats. The results of other methods are from their official reports. ”Direct” refers to the zero-shot setting.
Method HumanEval HumanEval-ET MBPP MBPP-ET BigCodeBench(full)
@@
AgentCoder [^1] Despite multiple attempts, we were unable to reproduce the reported AgentCoder results during our experiments. When contacted, the AgentCoder authors acknowledged they are revising their approach due to reproducibility challenges. We include their reported numbers for completeness but advise caution in direct comparisons.  (Huang et al., 2024b) 96.3 86.0 91.8 91.8
@@
AgentCoder [^1] Despite multiple attempts, we were unable to reproduce the reported AgentCoder results during our experiments. When contacted, the AgentCoder authors acknowledged they are revising their approach due to reproducibility challenges. We include their reported numbers for completeness but advise caution in direct comparisons.  (Huang et al., 2024b) 79.9 77.4 89.9 89.1
@@
Unlike multi-agent frameworks in Section 2 that depend heavily on LLM agents, PyCapsule achieves more efficient coordination with fewer agents, API calls, and reduced token usage.
In Table 1, PyCapsule achieves the state-of-the-art code generation tasks across HumanEval, HumanEval-ET, and BigCodeBench(full). While showing strong performance on MBPP, it falls short in comparison to AgentCoder and Qwen-2.5-Coder-Instruct which employs three-shot prompt engineering to stabilize code generation. Although it shows remarkable potential, we maintain a consistent prompting approach across all datasets to ensure comparability and fairness. To achieve this, we utilize our signature converter module to generate actionable function signatures. Remarkably, on HumanEval, the integration of Qwen-2.5-Coder-Instruct’s 7B model with PyCapsule achieves a 94.1% success rate compared to 92.7% by its 32B model at zero-shot. Furthermore, on BigCodeBench(full), PyCapsule with Qwen-2.5-Coder-Instruct 7B achieves a 25% improvement over the model’s standalone zero-shot performance of 41.0%, and exceeds the 32B variant (49.6%) by more than 15%. These substantial gains clearly demonstrate PyCapsule’s effectiveness in enhancing coding capabilities beyond what base LLMs can achieve independently, even when compared to much larger models.
Figure 2. Distribution of relative success ratios along the self-debugging attempts. The relative success ratio refers to the number of successful test cases with given attempts over the total number of successful test cases, measured by unit %. Results on all the datasets consistently show an exponential decreasing relative success ratio as the number of attempts increases.
Influence of Self-debugging Attempts:
To evaluate the influence of each self-debugging attempt on the success rate, we analyze the normalized and accumulated impacts of each incremental attempt using GPT-4-1106-preview, GPT-3.5-Turbo-1106, and Qwen-2.5-Coder-Instruct 7B. The success rate normalization was performed as follows: given $N$ problems, the number of successfully solved problems is denoted as $S_{0}$ while the rest $N_{1}=N-S_{0}$ unsolved problems will proceed to the next debugging attempt. The number of problems solved with $i\geq 1$ attempts is denoted as $S_{i}$, and $N_{i+1}=N_{i}-S_{i}$ represents the rest. The independent influence is defined as
@@
Figure 3. The plot illustrates the exponential decay in the normalized independent influence of self-debugging attempts across models. It shows the mean effectiveness of each debugging attempt, fitted to the exponential decay function $I(x)=a\cdot e^{-bx}$. The excellent fit of this function serves as an ideal guide for estimating the influence of successive attempts and can be used as a metric to compare the performance of different models.
This metric highlights the independent contribution of each attempt. We fitted the influence data to an exponential decay function and calculated the goodness of fit, obtaining $R^{2}$, which was close to one for every model and dataset, confirming an exponential decay pattern in debugging effectiveness. From this, we can clearly see the descending trend in the independent influence indicates diminishing effectiveness of debugging attempts, where problems requiring more attempts may need enriched conversation history or enhanced prompts to improve accuracy. The effectiveness approaches close to zero after five attempts for Qwen2.5-Coder-Instruct 7B, with GPT-4-Preview-1106 reaching this threshold at 3 attempts. Based on these findings, we established five debugging attempts as the maximum in our architecture.
@@ 4.3. Discussion
PyCapsule’s single-attempt accuracy varies considerably across datasets and model configurations, as shown in Figure 2. Our analysis reveals significant differences in first-attempt performance across frameworks, as detailed in Appendix B and Table 3. These performance differences highlight how framework design and prompt engineering significantly influence code generation capability even before self-debugging mechanisms are applied. Prior works (Wei et al., 2022; Chen et al., 2024; Patel et al., 2024) have demonstrated that prompt design critically affects model performance. PyCapsule leverages sophisticated system prompts(detailed in Appendix F) that are dynamically augmented in both code generation and fix modes, contributing to its strong first-attempt performance. When examining performance across multiple iterations, PyCapsule shows remarkable improvements, ultimately surpassing other self-debugging methods in overall success rates. This iterative improvement further validates our approach to error handling and feedback incorporation between agents.
Furthermore, an important advantage of PyCapsule is its reduced reliance on LLM-based agents, leading to reduced token consumption and improved computational efficiency. Existing frameworks like MapCoder can make up to 17 API calls with GPT models per problem on HumanEval and 12 calls on MBPP, following $k\times t$ debugging attempts, where $k$ is the number of retrieved exemplars and $t$ is the number of self-debugging attempts. In contrast, PyCapsule streamlines this process by limiting the programmer agent with a maximum of 5 attempts, requiring at most 6 LLM API calls for each problem because planning and code generation are handled within a single API call. Our empirical analysis shows that PyCapsule achieves superior code generation ability while maintaining efficient API usage, requiring an average of only 2.09, 1.38, and 1.55 API calls for HumanEval using GPT-3.5-Turbo-1106, GPT-4-Preview-1106, and Qwen-2.5-Coder-Instruct-7B, respectively. This optimization demonstrates the possibility of achieving high performance with significantly reduced computational overhead. For a detailed analysis of token usage and computational efficiency comparisons, see Appendix G
@@ Appendix A Contribution of Each Debugging Attempt
Appendix B First Attempt Accuracy
Table 3. Comparison of first-attempt accuracy across frameworks
Framework Model HumanEval HumanEval-ET MBPP MBPP-ET
MapCoder GPT-3.5 48.1 37.2 49.8 37.7
GPT-4 80.1 73.8 81.1 54.7
AgentCoder GPT-3.5 57.3 42.7 52.2 36.8
GPT-4 67.6 50.6 68.3 52.2
LDB GPT-3.5 73.8 - 67.6 -
Base LLM Qwen2.5-Coder-Instruct(7B) 88.4 84.1 83.5 71.7
PyCapsule GPT-3.5 81.9 $\pm$ 0.6 77.2 $\pm$ 1.5 74.6 $\pm$ 1.8 70.3 $\pm$ 1.3
GPT-4 92.0 $\pm$ 1.4 88.0 $\pm$ 3.7 70.6 $\pm$ 1.0 67.5 $\pm$ 0.8
Qwen2.5-Coder-7B 79.3 $\pm$ 1.4 81.3 $\pm$ 0.4 60.9 $\pm$ 0.7 59.7 $\pm$ 1.2
PyCapsule shows significantly better first-attempt performance compared to other frameworks like MapCoder and AgentCoder. Using GPT-3.5, PyCapsule achieves higher accuracy on benchmarks such as HumanEval (81.9% vs. 48.1% and 57.3%) and HumanEval-ET (77.2% vs. 37.2% and 42.7%). It also outperforms on MBPP and MBPP-ET benchmarks.
With GPT-4, PyCapsule remains competitive, though slightly behind MapCoder in some MBPP benchmarks. Compared to LDB with GPT-3.5, PyCapsule performs better on HumanEval and MBPP. However, Qwen2.5-Coder-Instruct(7B) outperforms PyCapsule in MBPP benchmarks due to its 3-shot prompting.
PyCapsule’s success is due to its carefully crafted prompts with clear instructions, contextual information, and encouragement of chain-of-thought reasoning. It also uses a well-defined programmer agent persona and optimized implementation details like temperature settings and token lengths.
Appendix C HumanEval Experiments
@@ Appendix D Pseudo-code of PyCapsule
Appendix E Supportive Modules (Detailed)
E.1. Signature Converter
  • The inferred function name.
  • A structured function signature (e.g., argument names and types).
  • An example function call derived from the test case.
### Required function name for your reference ‘foo()’
### Function signature for your reference - foo(arg_int: int)
### An example function call from private test cases - foo(4)
E.2. Example Call Detection
E.3. Error-Handling
Appendix F System Prompts
F.1. Code Generation Mode
@@ F.2. Fix Mode
Appendix G Token Usage
Table 4. Computational Resource Efficiency of PyCapsule: Token Usage and Debugging Attempt Analysis Across GPT-3.5, GPT-4, and Qwen2.5-Coder-Instruct Models on Standard Code Generation Benchmarks
Dataset Total Tokens Avg Tokens/Problem Avg Attempts/Problem
GPT-3.5-Turbo-1106
HumanEval 69,502.33 423.79 2.05
HumanEval-ET 75,263.67 458.92 2.10
MBPP 432,144.00 443.68 2.46
MBPP-ET 671,316.33 689.24 3.22
GPT-4-Preview-1106
HumanEval 79,342.50 483.80 1.35
HumanEval-ET 82,214.00 501.30 1.37
MBPP 779,440.33 800.25 2.03
MBPP-ET 1,044,909.33 1,072.80 2.79
Qwen2.5-Coder-Instruct
HumanEval 70,868.67 432.13 1.60
HumanEval-ET 74,151.33 452.14 1.61
MBPP 646,635.67 663.90 2.54
MBPP-ET 832,728.00 854.96 3.31
BigCodeBench 1,960,864.00 1,720.06 3.15
PyCapsule experiments employed a rigorous methodology, utilizing OpenAI’s ”tiktoken” library for GPT-3.5 and GPT-4 token counting, and Qwen2.5-Coder-Instruct’s native tokenizer for accurate measurement. We logged all inference calls, including debugging iterations, and calculated per-problem averages by dividing total token usage by the number of problems. Attempt counts were tracked by monitoring execution cycles until problems were solved or the maximum attempt limit was reached. It is essential to note that token counting methodologies vary across frameworks, making direct numerical comparisons challenging. However, PyCapsule’s significantly lower debugging attempt counts (1.35-3.31 across all models and datasets) highlight its efficiency in reaching solutions with minimal iterations.